OpenStudy (anonymous):

a buffer solution containing H2CO3 and NaHCO3 is to be prepared to maintain a pH of 7.00. What must be the ratio NaHO3/HCO in order to realize such as pH if Ka of carbonic acid is 4.3x10-7.

OpenStudy (hba):

@wio

OpenStudy (abmon98):

Henderson–Hasselbalch equation can be used to find the ph of a buffer solution or the ratio of conjugate acid and conjugate base pH = Pka + log (A-) / (HA) 7=4.3*10^-6+log(A-)/(HA)

OpenStudy (abb0t):

Then, take the inverse log to find the ratio and remove the log.